# ProxySPEX: Inference-Efficient Interpretability via Sparse Feature Interactions in LLMs

### v1

@@ Abstract
Large Language Models (LLMs) have achieved remarkable performance by capturing complex interactions between input features. To identify these interactions, most existing approaches require enumerating all possible combinations of features up to a given order, causing them to scale poorly with the number of inputs $n$. Recently, Kang et al. (2025) proposed SPEX, an information-theoretic approach that uses interaction sparsity to scale to $n\approx 10^{3}$ features. SPEX greatly improves upon prior methods but requires tens of thousands of model inferences, which can be prohibitive for large models. In this paper, we observe that LLM feature interactions are often hierarchical—higher-order interactions are accompanied by their lower-order subsets—which enables more efficient discovery. To exploit this hierarchy, we propose ProxySPEX, an interaction attribution algorithm that first fits gradient boosted trees to masked LLM outputs and then extracts the important interactions. Experiments across four challenging high-dimensional datasets show that ProxySPEX more faithfully reconstructs LLM outputs by 20% over marginal attribution approaches while using 10×10$\times$10 × fewer inferences than SPEX. By accounting for interactions, ProxySPEX identifies features that influence model output over 20% more than those selected by marginal approaches. Further, we apply ProxySPEX to two interpretability tasks. Data attribution, where we identify interactions among CIFAR-10 training samples that influence test predictions, and mechanistic interpretability, where we uncover interactions between attention heads, both within and across layers, on a question-answering task. ProxySPEX identifies interactions that enable more aggressive pruning of heads than marginal approaches.
@@ 1 Introduction
Figure 1: ProxySPEX requires $\sim$10×10$\times$10 × fewer inferences to achieve equally faithful explanations as SPEX for a sentiment classification and image-captioning task using a BERT and CLIP model respectively. LASSO faithfulness plateaus indicating limits of marginal approaches.
We show empirically that local (i.e., input specific) LLM feature interactions are often hierarchical: for an order $d$ interaction, an LLM includes lower-order interactions involving subsets of those $d$ features (see Figure 2). We use this to develop ProxySPEX, an interaction attribution algorithm that reduces the number of inferences compared to SPEX by 10×10$\times$10 × while achieving equally faithful explanations. ProxySPEX exploits this local hierarchical structure by first fitting gradient boosted trees (GBTs) as a proxy model to predict the output of LLMs on masked input sequences. Then, ProxySPEX extracts important interactions from the fitted GBTs [5].
@@
[…]
@@
    Feature identification. By accounting for interactions, ProxySPEX identifies influential features that impact model outputs more significantly than marginal approaches.
@@ 2 Related work and applications
Feature and interaction attribution. SHAP [1] and LIME [2] are widely used for model-agnostic feature attribution. SHAP uses the game-theoretic concept of Shapley values [9] for feature attribution, while LIME fits a sparse linear model [10]. Cohen-Wang et al. [11] also consider fitting a sparse linear model for feature attribution. Chen et al. [12] uses an information-theoretic approach for feature attributions. Other methods [13, 14] study model structure to derive feature attributions. Tsai et al. [3], Sundararajan et al. [15] and Bordt and von Luxburg [16] define extensions to Shapley values that consider interactions. Fumagalli et al. [17] provides a framework for computing several interaction attribution scores, but their approach does not scale past $n\approx 20$ features, which prevents them from being applied to modern ML problems that often consist of hundreds of features.
Fourier transforms and deep learning explainability. Several works theoretically study the spectral properties of transformers. Ren et al. [18] show transformers have sparse spectra and Hahn and Rofin [19], Abbe et al. [20] establish that they are low degree. Abbe et al. [21, 22] study the bias of networks learning interactions via a “staircase” property, i.e., using lower-order terms to learn high-order interactions. Sparsity and low degree structure is also empirically studied in [23, 24]. Kang et al. [25] shows that under sparsity in the Möbius basis [26], a representation closely related to Shapley values and the Fourier transform, interaction attributions can be computed efficiently. Kang et al. [4] use these insights to propose SPEX, the first robust interaction attribution algorithm to scale to the order of $n\approx 1000$ features. Gorji et al. [5] apply sparse Fourier transforms [27, 28, 29, 30] for computing Shapley values. They also provide an algorithm to extract the Fourier transform of tree-based models using a single forward pass.
@@ 3.1 Preliminaries
Value function. Let $\mathbf{x}$ be the input to the LLM consisting of $n$ features [^1] Features refer to inputs at a given granularity, e.g., tokens in an LLM or image patches in a vision model.. For $S\subseteq[n]$, where $[n]=1,\dotsc,n$, denote $\mathbf{x}_{S}$ as the masked input where we retain features indexed in $S$ and replace all others with the [MASK] token. For example, in the sentence $\mathbf{x}=$“The sequel truly elevated the original”, if $S=\{1,2,5,6\}$, $\mathbf{x}_{S}=$ “The sequel [MASK] [MASK] the original”. Masks can be more generally applied to any type of input such as image patches in a vision-language model. For a masked input $\mathbf{x}_{S}$ and LLM $f$, let $f(\mathbf{x}_{S})\in\mathbb{R}$ denote the output of the LLM under masking pattern $S$. The value function $f$ is problem dependent. For classification tasks, a common choice is the logit of the predicted class for unmasked input, $f(\mathbf{x})$. In generative tasks, $f(\mathbf{x}_{S})$ can represent the perplexity of generating the original output for the unmasked input. Since we focus on providing input-specific explanations, we suppress notation on $\mathbf{x}$ and denote $f(\mathbf{x}_{S})$ as $f(S)$.
@@
[…]
The parameters $F(T)$ are known as Fourier coefficients and capture the importance of an interaction of features in a subset $T$. Equation 1 represents an orthonormal transform onto a parity (XOR) basis [36]. For the rest of the paper, we use the terms Fourier coefficient and interaction interchangeably. Further, we refer to the set of Fourier coefficients $\{(T,F(T)):T\subseteq[n]\}$ as the spectrum.
@@
A faithful and sparse $\hat{f}$ allows straightforward computation of all popular feature or interaction attribution scores defined in the literature, e.g., Shapley, Banzhaf, Influence Scores, Faith-Shapley. Closed-form formulas for converting $F$ to various attribution indices are provided in Section A.1.
@@ 3.2 Empirical evidence of spectral hierarchies
For the top $k$ coefficients (i.e., interactions), DSR measures the average fraction of Fourier coefficients that exclude only one of the features $F(S\setminus\{i\})$. For example, an $f$ with $\mathcal{F}_{4}=\{\emptyset,\{1\},\{2\},\{1,3\}\}$ would have DSR of $\frac{1}{4}\left(1+1+1+\frac{1}{2}\right)=\frac{7}{8}$. High DSR implies that significant high-order interactions have corresponding significant lower-order Fourier coefficients. Figure 2 visualizes hierarchical interactions. Next, we show that two LLM based value functions have high DSR.
We take $20$ samples from a sentiment analysis task and an image captioning task [41]; see Section 4 for a detailed description and our choice of value function. We generate masks $S$ and apply SPEX until our learned value function has faithfulness ($R^{2}$) more than $0.9$. Figure 3 visualizes the DSR for various values of $k$, i.e., number of top interactions. DSR is consistently larger than $80\%$, indicating strong hierarchical structure. In Appendix B.2, we consider two additional metrics measuring hierarchical structure, and demonstrate that the top-$k$ interactions are faithful.
@@
Using GBTs to capture hierarchical Interactions. Tan et al. [42] proved that decision trees learn “staircase” functions, e.g., $f=x_{1}+x_{1}x_{2}+x_{1}x_{2}x_{3}$, effectively due to their greedy construction procedure. We empirically confirm this by comparing the performance of various proxy models on a synthetic hierarchical function (i.e., sum of staircase functions resembling Figure 2) as well as the Sentiment dataset in Appendix Figure 12. Section B.4 details the simulation set-up. GBTs vastly outperform other proxy models, indicating their natural ability to identify hierarchical interactions with limited training data. Interestingly, GBTs outperform random forests as well. This is because random forests are ineffective at learning hierarchical functions [43], i.e., sums of staircases, while GBT-like algorithms disentangle sums effectively [44].
@@ 3.3 ProxySPEX via Gradient Boosted Trees to fit hierarchies
The ProxySPEX algorithm (see Figure 4):
@@
Step 3 - Fourier extraction. We use Gorji et al. [5] to extract the Fourier representation of the fitted GBTs in a single forward pass; see Section A.2. With $T$ trees of depth $d$ there are at most $O(T4^{d})$ non-zero Fourier coefficients. To improve interpretability, we sparsify the extracted representation by keeping only the top $k$ Fourier coefficients. Fig. 5 shows that only $\approx 200$ Fourier coefficients are needed to achieve equivalent faithfulness for a sentiment classification and image captioning (MS-COCO) dataset. Additional results regarding the sparsity of Fourier spectra learned by GBTs are in Appendix B.3.
Step 4 (Optional): Coefficient refinement via regression. Optionally we regress the extracted, sparsified Fourier coefficients on the collected data $\mathcal{D}$ to improve the estimation. This step is included if it leads to lower CV error.
@@ 4 Results
Baselines and hyperparameters. For marginal feature attributions, we use the LASSO. We use the same datasets at [4] and add MS-COCO for an additional modality. It was shown in [4] that popular marginal metrics such as SHAP are significantly less faithful than the LASSO, e.g., have $R^{2}<0$. We use the LASSO implementation from scikit-learn, and choose the $l_{1}$ regularization parameter via $5$-fold CV. For interaction indices, we compare ProxySPEX to SPEX. Due to the scale of $n$ in our experiments, we cannot compare methods for computing interaction indices such as Faith-Shapley, Faith-Banzhaf, and Shapley-Taylor using SHAP-IQ [17], and SVARM-IQ [50], because they enumerate all possible interactions, making them computationally infeasible. For ProxySPEX, a list of GBT hyper-parameters we tune over are in Appendix B.
@@ 4.1 Faithfulness
Figure 6 shows average faithfulness over 1,000 test masks per sample. ProxySPEX outperforms LASSO with limited inferences and continues to improve where LASSO plateaus, indicating that it is learning influential interactions. While SPEX is often faster for the same number of masks, SPEX needs additional inference time to match $R^{2}$, making ProxySPEX faster overall. For the smaller DistilBERT model under the sentiment analysis task, the wall clock speedup is $\sim$3×3$\times$3 ×, while with the bigger CLIP-ViT-B/32 model with MS-COCO we see $\sim$5×5$\times$5 × speedup (See Section B.5).
@@ 4.2 Feature Identification
Solving Eq. 4 for an arbitrary $\hat{f}$ presents a challenging combinatorial optimization problem. However, ProxySPEX and SPEX represent $\hat{f}$ as a sparse Fourier transform. This representation facilitates solving the optimization as a tractable linear integer program. The sparsity of the extracted Fourier representation ensures that the time required to solve this program is negligible compared to sampling the LLM and fitting the GBTs. Full details of the construction of this program are given in Section A.3. Under LASSO, Eq. 4 is easily solved through selecting features by the size of their coefficients. We measure the removal ability of different attribution methods when we collect $8n\log_{2}(n)$ training masks and plot the result in Figure 7. By accounting for interactions, ProxySPEX identifies significantly more influential features than the LASSO. Apart from the sentiment analysis task, SPEX does not collect enough training masks to outperform the LASSO.
@@ 5.1 Data Attribution via Non-Linear Datamodels
$f(S)\triangleq(\text{logit for }c\text{ on }\mathbf{z})-(\text{highest incorrect logit on }\mathbf{z}),\;\;\text{when }g_{\theta}\text{ is trained on}S.$ (5)
@@
Results. We visualize interactions learned by ProxySPEX in Figure 8 for randomly selected CIFAR-10 test points. Experimental details are in Section C.1. ProxySPEX identifies highly similar training samples (redundancies) as well as synergistic interactions between samples of different classes. See Appendix C.1 for examples of other randomly selected test samples.

### v2

@@ Abstract
Large Language Models (LLMs) have achieved remarkable performance by capturing complex interactions between input features. To identify these interactions, most existing approaches require enumerating all possible combinations of features up to a given order, causing them to scale poorly with the number of inputs $n$. Recently, Kang et al. (2025) proposed SPEX, an information-theoretic approach that uses interaction sparsity to scale to $n\approx 10^{3}$ features. SPEX greatly improves upon prior methods but requires tens of thousands of model inferences, which can be prohibitive for large models. In this paper, we observe that LLM feature interactions are often hierarchical—higher-order interactions are accompanied by their lower-order subsets—which enables more efficient discovery. To exploit this hierarchy, we propose ProxySPEX, an interaction attribution algorithm that first fits gradient boosted trees to masked LLM outputs and then extracts the important interactions. Experiments across four challenging high-dimensional datasets show that ProxySPEX more faithfully reconstructs LLM outputs by 20% over marginal attribution approaches while using $10\times$ fewer inferences than SPEX. By accounting for interactions, ProxySPEX efficiently identifies the most influential features, providing a scalable approximation of their Shapley values. Further, we apply ProxySPEX to two interpretability tasks. Data attribution, where we identify interactions among CIFAR-10 training samples that influence test predictions, and mechanistic interpretability, where we uncover interactions between attention heads, both within and across layers, on a question-answering task. The ProxySPEX algorithm is available at https://github.com/mmschlk/shapiq.
@@ 1 Introduction
Figure 1: ProxySPEX requires $\sim$$10\times$ fewer inferences to achieve equally faithful explanations as SPEX for a sentiment classification and image-captioning task using a BERT and CLIP model respectively. LASSO faithfulness plateaus indicating limits of marginal approaches.
We show empirically that local (i.e., input specific) LLM feature interactions are often hierarchical: for an order $d$ interaction, an LLM includes lower-order interactions involving subsets of those $d$ features (see Figure˜2). We use this to develop ProxySPEX, an interaction attribution algorithm that reduces the number of inferences compared to SPEX by $10\times$ while achieving equally faithful explanations. ProxySPEX exploits this local hierarchical structure by first fitting gradient boosted trees (GBTs) as a proxy model to predict the output of LLMs on masked input sequences. Then, ProxySPEX extracts important interactions from the fitted GBTs [5].
@@
    Faithfulness. ProxySPEX learns more faithful representations of LLM outputs than marginal approaches $(\approx$$15\%\text{ to }25\%)$ on average across datasets as we vary the number of inferences. Figure˜1 compares explanation faithfulness of ProxySPEX to marginal attributions and SPEX.
@@
    Feature identification. By accounting for interactions, ProxySPEX identifies influential features that impact model outputs more significantly than marginal approaches, and can approximate Shapley values better than KernalSHAP in the low-inference regime.
@@ 2 Related work and applications
Feature and interaction attribution. SHAP [1] and LIME [2] are widely used for model-agnostic feature attribution. SHAP uses the game-theoretic concept of Shapley values [9] for feature attribution, while LIME fits a sparse linear model [10]. Cohen-Wang et al. [11] also consider fitting a sparse linear model for feature attribution. Chen et al. [12] uses an information-theoretic approach for feature attributions. Other methods [13, 14] study model structure to derive feature attributions. Sundararajan et al. [15] and Bordt and von Luxburg [16] define extensions to Shapley values that consider interactions. Fumagalli et al. [17] provides a framework for computing several interaction attribution scores, but their approach does not scale past $n\approx 20$ features, which prevents them from being applied to modern ML problems that often consist of hundreds of features. Note that some feature attribution approaches such as LIME and Faith SHAP [3] are formulated explicitly as a function approximation, while others are defined axiomatically such as SHAP, though one can typically construct equivalent function approximation objectives with a suitable distance metric.
Fourier transforms and deep learning explainability. Several works theoretically study the spectral properties of transformers. Ren et al. [18] show transformers have sparse spectra and Hahn and Rofin [19], Abbe et al. [20] establish that they are low degree. Abbe et al. [21, 22] study the bias of networks learning interactions via a “staircase” property, i.e., using lower-order terms to learn high-order interactions. Sparsity and low degree structure is also empirically studied in [23, 24]. Kang et al. [25] shows that under sparsity in the Möbius basis [26], a representation closely related to Shapley values and the Fourier transform, interaction attributions can be computed efficiently. Mohammadi et al. [27] also learn a sparse Möbius representation for computing Shapley values. Kang et al. [4] use these insights to propose SPEX, the first robust interaction attribution algorithm to scale to the order of $n\approx 1000$ features. Gorji et al. [5] apply sparse Fourier transforms [28, 29, 30, 31] for computing Shapley values. They also provide an algorithm to extract the Fourier transform of tree-based models using a single forward pass.
SPEX. We refer to the algorithm proposed in this manuscript as ProxySPEX, in reference to SPEX, since both works exploit a sparse interaction prior to reduce computational and sample budget. SPEX uses an algebraic structured sampling scheme, coupled with error correction decoding procedures to efficiently compute the interactions in the form of a Fourier transform. In contrast, ProxySPEX uses random samples to learn a proxy model that implicitly exploits the sparse interaction priors and our newly proposed hierarchical prior.
@@ 3.1 Preliminaries
Value function. Let $\mathbf{x}$ be the input to the LLM consisting of $n$ features [^1] Features refer to inputs at a given granularity, e.g., tokens in an LLM or image patches in a vision model.. For $S\subseteq[n]$, where $n]=1,\dotsc,n$, denote $\mathbf{x}_{S}$ as the masked input where we retain features indexed in $S$ and replace all others with the [MASK] token. For example, in the sentence $\mathbf{x}=$“The sequel truly elevated the original”, if $S=\{1,2,5,6\}$, $\mathbf{x}_{S}=$ “The sequel [MASK] [MASK] the original”. Masks can be more generally applied to any type of input such as image patches in a vision-language model. For a masked input $\mathbf{x}_{S}$ and LLM $f$, let $f(\mathbf{x}_{S})\in\mathbb{R}$ denote the output of the LLM under masking pattern $S$. The value function $f$ is problem dependent. For classification tasks, a common choice is the logit of the predicted class for unmasked input, $f(\mathbf{x})$. In generative tasks, $f(\mathbf{x}_{S})$ can represent the perplexity of generating the original output for the unmasked input. Since we focus on providing input-specific explanations, we suppress notation on $\mathbf{x}$ and denote $f(\mathbf{x}_{S})$ as $f(S)$.
@@
$\text{Transform: }F(T)=\frac{1}{2^{n}}\sum_{S\subseteq[n]}(-1)^{|S\cap T|}f(S),\qquad\text{Inverse: }f(S)=\sum_{T\subseteq[n]}(-1)^{|T\cap S|}F(T).$ (1)
The parameters $F(T)$ are known as Fourier coefficients and capture the importance of an interaction of features in a subset $T$. Equation˜1 represents an orthonormal transform onto a parity (XOR) basis [37]. For the rest of the paper, we use the terms Fourier coefficient and interaction interchangeably. Further, we refer to the set of Fourier coefficients $\{(T,F(T)):T\subseteq[n]\}$ as the spectrum.
@@
A faithful and sparse $\hat{f}$ allows straightforward computation of all popular feature or interaction attribution scores defined in the literature, e.g., Shapley, Banzhaf, Influence Scores, Faith-Shapley. Closed-form formulas for converting $F$ to various attribution indices are provided in Section˜A.1.
@@ 3.2 Empirical evidence of spectral hierarchies
For the top $k$ coefficients (i.e., interactions), DSR measures the average fraction of Fourier coefficients that exclude only one of the features $F(S\setminus\{i\})$. For example, an $f$ with $\mathcal{F}_{4}=\{\emptyset,\{1\},\{2\},\{1,3\}\}$ would have DSR of $\frac{1}{4}\left(1+1+1+\frac{1}{2}\right)=\frac{7}{8}$. High DSR implies that significant high-order interactions have corresponding significant lower-order Fourier coefficients. Figure˜2 visualizes hierarchical interactions. Next, we show that two LLM based value functions have high DSR.
We take $20$ samples from a sentiment analysis task and an image captioning task [42]; see Section˜4 for a detailed description and our choice of value function. We generate masks $S$ and apply SPEX until our learned value function has faithfulness ($R^{2}$) more than $0.9$. Figure˜3 visualizes the DSR for various values of $k$, i.e., number of top interactions. DSR is consistently larger than $80\%$, indicating strong hierarchical structure. In Appendix B.2, we consider two additional metrics measuring hierarchical structure, and demonstrate that the top-$k$ interactions are faithful.
@@
Using GBTs to capture hierarchical Interactions. Tan et al. [43] proved that decision trees learn “staircase” functions, e.g., $f=x_{1}+x_{1}x_{2}+x_{1}x_{2}x_{3}$, effectively due to their greedy construction procedure. We empirically confirm this by comparing the performance of various proxy models on a synthetic hierarchical function (i.e., sum of staircase functions resembling Figure˜2) as well as the Sentiment dataset in Appendix Figure˜13. Section˜B.4 details the simulation set-up. GBTs vastly outperform other proxy models, indicating their natural ability to identify hierarchical interactions with limited training data. Interestingly, GBTs outperform random forests as well. This is because random forests are ineffective at learning hierarchical functions [44], i.e., sums of staircases, while GBT-like algorithms disentangle sums effectively [45].
@@ 3.3 ProxySPEX via Gradient Boosted Trees to fit hierarchies
The ProxySPEX algorithm (see Figure˜4):
@@
Step 3 - Fourier extraction. We use Gorji et al. [5] to extract the Fourier representation of the fitted GBTs in a single forward pass; see Section˜A.2. With $T$ trees of depth $d$ there are at most $O(T4^{d})$ non-zero Fourier coefficients [5]. To improve interpretability, we sparsify the extracted representation by keeping only the top $k$ Fourier coefficients. Fig. 5 shows that only $\approx 200$ Fourier coefficients are needed to achieve equivalent faithfulness for a sentiment classification and image captioning (MS-COCO) dataset. Additional results regarding the sparsity of Fourier spectra learned by GBTs are in Appendix B.3.
Step 4 (Optional): Coefficient refinement via regression. As a final step, we optionally regress the extracted, top $k$ Fourier coefficients on the collected data $\mathcal{D}$ to improve the estimation. Empirically we observe this step is can sometimes marginally improve performance, but seldom negatively impacts performance. This step is included if it leads to lower CV error.
@@ 4 Results
Baselines and hyperparameters. For marginal feature attributions, we use the LASSO. We use the same datasets at [4] and add MS-COCO for an additional modality. It was shown in [4] that popular marginal metrics such as SHAP are significantly less faithful than the LASSO, e.g., have $R^{2}<0$. We use the LASSO implementation from scikit-learn, and choose the $l_{1}$ regularization parameter via $5$-fold CV. For interaction indices, we compare ProxySPEX to SPEX. Due to the scale of $n$ in our experiments, we cannot compare methods for computing interaction indices such as Faith-Shapley, Faith-Banzhaf, and Shapley-Taylor using SHAP-IQ [17], and SVARM-IQ [51], because they enumerate all possible interactions, making them computationally infeasible. For ProxySPEX, a list of GBT hyper-parameters we tune over are in Appendix˜B.
@@ 4.1 Faithfulness
Figure˜6 shows average faithfulness over 1,000 test masks per sample. ProxySPEX outperforms LASSO with limited inferences and continues to improve where LASSO plateaus, indicating that it is learning influential interactions. While SPEX is often faster for the same number of masks, SPEX needs additional inference time to match $R^{2}$, making ProxySPEX faster overall. For the smaller DistilBERT model under the sentiment analysis task, the wall clock speedup is $\sim$$3\times$, while with the bigger CLIP-ViT-B/32 model with MS-COCO we see $\sim$$5\times$ speedup (See Section˜B.6).
@@ 4.2 Feature Identification
Solving Eq. 4 for an arbitrary $\hat{f}$ presents a challenging combinatorial optimization problem. However, ProxySPEX and SPEX represent $\hat{f}$ as a sparse Fourier transform. This representation facilitates solving the optimization as a tractable linear integer program. The sparsity of the extracted Fourier representation ensures that the time required to solve this program is negligible compared to sampling the LLM and fitting the GBTs. Full details of the construction of this program are given in Section˜A.3. Under LASSO, Eq. 4 is easily solved through selecting features by the size of their coefficients. We measure the removal ability of different attribution methods when we collect $8n\log_{2}(n)$ training masks and plot the result in Figure˜7. By accounting for interactions, ProxySPEX identifies significantly more influential features than the LASSO. Apart from the sentiment analysis task, SPEX does not collect enough training masks to outperform the LASSO.
4.3 Shapley Value Approximation
We evaluate the performance of ProxySPEX for efficiently approximating Shapley values. Across all tasks, we first run KernelSHAP with 10,000 test masks and treat these approximated Shapley values as ground truth. We measure the recall of the top ten highest-magnitude Shapley values for KernelSHAP and ProxySPEX under $\alpha\cdot n\log_{2}(n)$ inferences with multipliers $\alpha\in\{0.25,0.5,0.75,1.0\}$. For this inference budget, competing algorithms such as LeverageSHAP [53] and SVARM [54] struggle to provide accurate approximations. We find ProxySPEX initially provides a better coarse approximation than KernelSHAP (Figure 14). However, since ProxySPEX is optimized for faithfulness and does not rely on the Shapley kernel, it is eventually surpassed by KernelSHAP with enough inferences. Additional results under mean squared error are included in Appendix B.5.
Figure 8: Recall of the top ten Shapley values after $\alpha\cdot n\log_{2}(n)$ inferences for multipliers $\alpha\in\{0.25,0.5,0.75,1.0\}$. For small $\alpha$, ProxySPEX is superior at recovering the most significant features, while KernalSHAP outperforms as $\alpha$ increases. Error bands indicate the standard deviation across ten different runs of the algorithms.
@@ 5.1 Data Attribution via Non-Linear Datamodels
[…]
@@
Results. We visualize interactions learned by ProxySPEX in Figure˜9 for randomly selected CIFAR-10 test points. Experimental details are in Section˜C.1. ProxySPEX identifies highly similar training samples (redundancies) as well as synergistic interactions between samples of different classes. See Appendix C.1 for examples of other randomly selected test samples.